### v1

@@ 2.2 $\Gamma$-ideals, radicals, primes
Equivalently, $I$ is an ideal of the underlying semiring $(T,+,\cdot,0,1)$ (since $u_{\gamma}$ is a unit for every $\gamma$).
@@
For a $\Gamma$-ideal $I$, define its $\Gamma$-radical by
@@
For any $\Gamma$-ideal $I$, the set $\sqrt[\Gamma]{\,\cdot\,}(I)$ is a $\Gamma$-ideal containing $I$.
@@
If $x^{n}\in I$ then clearly $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$; hence $I\subseteq\sqrt[\Gamma]{\,\cdot\,}(I)$. Closure under addition and multiplication by arbitrary elements is standard: if $x^{m}\in I$ and $y^{n}\in I$, then $(x+y)^{m+n}$ expands into a sum of monomials each divisible by $x^{m}$ or $y^{n}$, hence lies in $I$; similarly $(t\cdot x)^{m}=t^{m}x^{m}\in I$. Since $\Gamma$-ideal closure is equivalent to ideal closure in our setting (Definition 2.4), we are done.
@@
Let $I$ be a $\Gamma$-ideal. Then
@@
($\subseteq$) If $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$, pick $n$ with $x^{n}\in I$. If $P$ is prime and contains $I$, then $x^{n}\in P$ implies $x\in P$. So $x$ lies in every prime containing $I$.
@@
Finally, $I\subseteq P$ by construction, and $x\notin P$ because $x\in P$ would force $x^{1}\in P$, contradicting $P\in\mathcal{S}$. Thus $x$ is not in the intersection of primes containing $I$.
@@ 2.3 The $\Gamma$-spectrum and $\Gamma$-Zariski topology
For a $\Gamma$-ideal $I$, define
@@ 2.4 The standard-cover equivalence lemma
($\Rightarrow$) Let $I=\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$. The equality of opens says: for every prime $P$, if $f\notin P$ then $f_{i}\notin P$ for some $i$. Equivalently, if $f_{i}\in P$ for all $i$ (i.e. $I\subseteq P$), then $f\in P$. Thus $f$ lies in the intersection of all primes containing $I$, hence $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ by Lemma 2.11.
($\Leftarrow$) Suppose $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$. Let $P\in D_{\Gamma}(f)$, so $f\notin P$. If $P$ contained $I$, then $f\in\sqrt[\Gamma]{\,\cdot\,}(I)\subseteq\sqrt[\Gamma]{\,\cdot\,}(P)=P$, contradicting $f\notin P$. Hence $I\not\subseteq P$, so some generator $f_{i}\notin P$, i.e. $P\in D_{\Gamma}(f_{i})$. Thus $D_{\Gamma}(f)\subseteq\bigcup_{i}D_{\Gamma}(f_{i})$, and the reverse inclusion holds because each $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ is automatic from the displayed radical inclusion: $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ implies $D_{\Gamma}(f)\supseteq D_{\Gamma}(f_{i})$ for each $i$.
@@
From Theorem 2.19 we have $f\in\sqrt[\Gamma]{\,\cdot\,}(\langle f_{1},\dots,f_{n}\rangle_{\Gamma})$, so $f^{N}\in\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$ for some $N$. By definition of generated ideal, $f^{N}$ is a finite sum of multiples of the generators, giving the claim.
@@ 2.6 Sheaf condition on the basis
    (existence of gluing) For any sections $s_{i}\in T_{f_{i}}$ such that their restrictions agree on each overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{j})=D_{\Gamma}(f_{i}f_{j})$, there exists $s\in T_{f}$ with $s|_{D_{\Gamma}(f_{i})}=s_{i}$ for all $i$.
@@
because equality in localization is detected by multiplication by a power of the localized element. Let $m:=\max_{i}m_{i}$. Then $f_{i}^{m}p=f_{i}^{m}q$ for every $i$. Multiply the relation (2.1) by itself $m$ times to obtain a decomposition
@@ 3.3 Affine anti-equivalence
contravariantly functorial in both variables. Moreover, for each $T$ the unit of this equivalence is the canonical isomorphism of Theorem 3.3.
@@ 4.1 Triadic bracket on the structure sheaf
Fix the $\Gamma$-semiring units $u_{\gamma}$ as in Definition 2.2.
@@ 4.2 A Nambu/Filippov-type identity
where the right-hand side uses the addition of $A$.

### v2

@@ 2.2 $\Gamma$-ideals, radicals, primes
Equivalently, II is an ideal of the underlying semiring $(T,+,\cdot,0,1)$ (since $u_{\gamma}$ is a unit for every $\gamma$).
@@
For a $\Gamma$-ideal II, define its $\Gamma$-radical by
@@
For any $\Gamma$-ideal II, the set $\sqrt[\Gamma]{\,\cdot\,}(I)$ is a $\Gamma$-ideal containing II.
@@
If $x^{n}\in I$ then clearly $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$; hence $I\subseteq\sqrt[\Gamma]{\,\cdot\,}(I)$. Closure under addition and multiplication by arbitrary elements is standard: if $x^{m}\in I$ and $y^{n}\in I$, then $(x+y)^{m+n}$ expands into a sum of monomials each divisible by $x^{m}$ or $y^{n}$, hence lies in II; similarly $(t\cdot x)^{m}=t^{m}x^{m}\in I$. Since $\Gamma$-ideal closure is equivalent to ideal closure in our setting (Definition 2.4), we are done.
@@
Let II be a $\Gamma$-ideal. Then
@@
($\subseteq$) If $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$, pick $n$ with $x^{n}\in I$. If $P$ is prime and contains II, then $x^{n}\in P$ implies $x\in P$. So $x$ lies in every prime containing II.
@@
Finally, $I\subseteq P$ by construction, and $x\notin P$ because $x\in P$ would force $x^{1}\in P$, contradicting $P\in\mathcal{S}$. Thus $x$ is not in the intersection of primes containing II.
@@ 2.3 The $\Gamma$-spectrum and $\Gamma$-Zariski topology
For a $\Gamma$-ideal II, define
@@ 2.4 The standard-cover equivalence lemma
($\Rightarrow$) Let $I=\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$. The equality of opens says: for every prime $P$, if $f\notin P$ then $f_{i}\notin P$ for some ii. Equivalently, if $f_{i}\in P$ for all $i$ (i.e. $I\subseteq P$), then $f\in P$. Thus $f$ lies in the intersection of all primes containing II, hence $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ by Lemma 2.11.
($\Leftarrow$) Suppose $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$. Let $P\in D_{\Gamma}(f)$, so $f\notin P$. If $P$ contained II, then $f\in\sqrt[\Gamma]{\,\cdot\,}(I)\subseteq\sqrt[\Gamma]{\,\cdot\,}(P)=P$, contradicting $f\notin P$. Hence $I\not\subseteq P$, so some generator $f_{i}\notin P$, i.e. $P\in D_{\Gamma}(f_{i})$. Thus $D_{\Gamma}(f)\subseteq\bigcup_{i}D_{\Gamma}(f_{i})$, and the reverse inclusion holds because each $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ is automatic from the displayed radical inclusion: $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ implies $D_{\Gamma}(f)\supseteq D_{\Gamma}(f_{i})$ for each ii.
@@
From Theorem 2.19 we have $f\in\sqrt[\Gamma]{\,\cdot\,}(\langle f_{1},\dots,f_{n}\rangle_{\Gamma})$, so $f^{N}\in\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$ for some NN. By definition of generated ideal, $f^{N}$ is a finite sum of multiples of the generators, giving the claim.
@@ 2.6 Sheaf condition on the basis
    (existence of gluing) For any sections $s_{i}\in T_{f_{i}}$ such that their restrictions agree on each overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{j})=D_{\Gamma}(f_{i}f_{j})$, there exists $s\in T_{f}$ with $s|_{D_{\Gamma}(f_{i})}=s_{i}$ for all ii.
@@
because equality in localization is detected by multiplication by a power of the localized element. Let $m:=\max_{i}m_{i}$. Then $f_{i}^{m}p=f_{i}^{m}q$ for every ii. Multiply the relation (2.1) by itself $m$ times to obtain a decomposition
@@ 3.3 Affine anti-equivalence
contravariantly functorial in both variables. Moreover, for each $T$ the unit of this equivalence is the canonical isomorphism of Theorem $3.$
@@ 4.1 Triadic bracket on the structure sheaf
Fix the $\Gamma$-semiring units $u_{\gamma}$ as in Definition $2.$
@@ 4.2 A Nambu/Filippov-type identity
where the right-hand side uses the addition of AA.